Box has surface area 24. Its length, width, height are doubled; new surface area = N
Answer by Alan3354(21584)

Box has surface area 24. Its length, width, height are doubled; new surface area = N
-------------
Area is a function of the square of the dimensions. The dims are 2x, so the area is 4x, = 96.
----------
To do it in detail:
Area = 2(l*h + l*w + w*h) length, width and height
New area = 2(2l*2h + 2l*2w + 2w*2h)
= 2(4l*h + 4l*w + 4w*h)
= 8(l*h + l*w + w*h) which is 4 times the original 24.